# Tawa: Automatic Warp Specialization for Modern GPUs with Asynchronous References

### v1

@@ I Introduction
To address these challenges, we propose Tawa 111A shorthand for Task-Aware Warp Specialization with Aref. The implementation has been upstreamed to OpenAI Triton: https://github.com/triton-lang/triton/tree/aref_auto_ws, an automated compiler that systematically analyzes computation graphs generated from a tile-based DSL, partitions programs into warp-specialized components, and manages both communication and pipelining. Our contributions are as follows:
We propose asynchronous references (aref), an abstraction for warp-level communication. aref expresses asynchronous data transfer intent without exposing low-level hardware primitives, providing a clean and portable representation at the IR level.
We present Tawa, an automatic compilation flow for warp specialization on NVIDIA GPUs. Tawa takes in a Triton program, performs task-aware partitioning across warp groups, applies multi-granularity pipelining, and generates high-performance PTX code for GPU execution.
@@ II-B Existing GPU Programming Models
To further raise the abstraction level, domain-specific compiler infrastructures have been proposed. TVM [14] pioneered an end-to-end compilation framework with automated scheduling [15, 16], while Relax [17] extends it with a graph-level IR for dynamic workloads. Fireiron [18] also leverages a schedule language for specifying the tiling and data movement, and Graphene [19] further enhances this idea by providing an IR for mapping tensors with various layouts to Tensor Core instructions. While powerful, these systems were developed prior to Hopper and do not directly exploit TMA and hardware warp specialization, leaving significant performance untapped. Similarly, lots of efforts on GPU warp specialization [20, 21, 22, 23] were proposed before the Hopper architecture, which do not effectively use the latest TMA features.
Motivated by the need for finer control, a new class of tile-based kernel languages has emerged. Triton [8], built on MLIR [24], allows programmers to write efficient kernels in Python while abstracting tiling and memory management. Despite its success in democratizing kernel development, Triton still lacks inherent abstractions for warp specialization, leaving a performance gap on the Hopper architecture. Recent work such as TileLang [25] and Mosaic [26] extends this direction by introducing explicit APIs for dataflow scheduling and pipeline composition. Cypress[27] proposes a task-based programming model that abstracts complex data movement and synchronization through logical computation descriptions and mapping specifications. However, programmers are still required to manually manage the multi-level hierarchy of threads, warps, and SMs, along with the corresponding data decomposition. Gluon [28] proposes an IR for Triton that adds layout abstractions but still requires programmers to explicitly incorporate communication operations. These approaches provide expert users with expressive control, but they continue to impose significant responsibility for synchronization and resource management, which limits programmer productivity.
To this end, we argue that users should not have to cope with low-level hardware details. Instead, warp-specialized programs should be generated automatically. Tawa bridges these gaps by introducing aref as an IR-level abstraction for warp communication and automatically generating warp-specialized, TMA-enabled pipelines, thereby achieving high performance without imposing low-level programming burden.
@@ III-A Overview
Fig. 2(a) illustrates the overall design of Tawa. Since Triton [8] has become the de facto backend for PyTorch [29, 30] and is widely adopted for writing efficient GPU kernels, we select Triton as the frontend interface of our compiler. Programmers write kernels in Triton-Python with tiled computation and TMA communication in a CTA (Fig. 2(b)), which are first translated into the standard Triton-MLIR representation. Building upon this infrastructure, Tawa introduces a series of compiler passes that automate the generation of warp-specialized programs directly from high-level Python code.
@@ III-C1 Partition Annotation
We perform a backward traversal along the use-def chains starting at the kernel’s side-effecting sinks (e.g., Store). During the traversal, we attach a semantic tag to every node by inspecting its effects. Nodes that contribute to address computation, such as the pointer arithmetic for data transfers, are marked as iteration statements (orange edges). Nodes that transform or consume a tile for actual computation, such as WGMMA, are marked as tile statements (blue edges). Importantly, iteration statements are not necessarily contiguous in the IR, as highlighted in Fig. 2(b).
A typical division on Hopper is to create a load WG as the producer and a compute WG as the consumer, as shown in Fig. 4. Tawa annotates the iteration statements and their dominated TMALoads in the producer partition (WG0), and the remaining tile statements in the consumer partition (WG1). Notice the epilogue is also attached to WG1 to make sure it is actually executed by some WGs. For Blackwell and later generations of GPUs that support more WGs, Tawa can also partition the IR based on available warp roles on GPUs by annotating the semantic roles during graph traversal.
@@ III-C2 Loop Distribution
After the operations are annotated, we need to construct the actual IR that divides the workload for each WG. For each cross-partition edge (e.g., TMALoadLocalAlloc in Fig. 5(a)), we create an aref tensor with size , denoting a -slot cyclic buffer. A put operation is created on the producer side that places each TMALoad result into an aref slot. The slot index is computed as the iteration modulo the buffer size , allowing the buffer slots to be reused (L15, L23 in Fig. 2(c)). Similarly, a get operation is created on the consumer side that gets the data from the same aref slot. After the data are used for computation in WGMMA, an additional consumed operation is inserted at the end to indicate the slot has been freed. As mentioned in § III-B, aref supports grouping multiple values in the buffer to share the mbarrier. Therefore, a straightforward optimization when creating aref is to check whether the loaded tensors are used in the same WGMMA. If so, we can put a tuple of tensors in the aref buffer and share the channel for communication between WGs. The resulting computation graph is shown in Fig. 5(b). Since aref has already carried buffers, the LocalAlloc operations can also be eliminated but to directly read from the aref buffer.
@@ V-A Experimental Setup
Tawa is built on top of Triton [8] (commit: 0c7edf) with about 4K lines of C++ and introduces a compact MLIR dialect that encapsulates aref operations. We compare Tawa against a set of state-of-the-art GPU programming frameworks and libraries, including the closed-source commercial library cuBLAS [39] (v12.7), the NVIDIA open-source CUDA linear algebra library CUTLASS [9] (v4.0.0), and the baseline Triton under the same commit. We also include two academic frameworks: ThunderKittens[10] (6c27e2), a CUDA-based AI kernel library, and TileLang[25] (bcfc83), a TVM [14]-based DSL. To ensure a fair comparison, all baselines are hand-tuned with a fixed set of tile sizes (combinations of ). In the default Triton implementation, the official software pipelining is enabled, with pipeline depth predetermined to balance overlap between communication and computation. For kernels generated by Tawa, the size of the aref and the depth of the MMA pipeline are selected manually to maximize performance, providing each approach with equivalent opportunity to exploit the underlying hardware.
@@ V-B Matrix Multiplication
We evaluate GEMM kernels with matrix dimensions of while sweeping the inner dimension from to , and report the throughput for FP16 and FP8. As shown in Fig. 8, across both precisions, Tawa achieves the same level of performance as the highly optimized kernel library cuBLAS while outperforming the other general-purpose frameworks on most shapes. Tawa reaches up to 79% hardware utilization in both FP16 and FP8.
In FP16, Tawa delivers average speedups of 1.01, 1.13, 1.23, and 1.09 over cuBLAS, Triton, TileLang, and ThunderKittens, respectively. The improvements over Triton highlight the advantage of overlapping TMA-driven data movement with WGMMA execution. In contrast, Triton employs an Ampere-style software pipelining scheme for asynchronous copies, which is less effective on Hopper than the hardware-backed warp-specialization strategy in Tawa. TileLang and ThunderKittens are extensively tuned for large , leading to stronger performance than Tawa when .
In FP8, the benefits of automatic warp specialization are more pronounced, particularly at small . On average, Tawa achieves 1.06, 1.02, 2.40, and 1.24 speedup over cuBLAS, Triton, TileLang, and ThunderKittens, respectively. TileLang encounters layout-management challenges for FP8 WGMMA, yielding an inferior implementation that lags other frameworks; at , Tawa is 3.99 faster than TileLang. ThunderKittens appears less carefully tuned for FP8, resulting in up to 1.61 lower throughput at small . Overall, these results indicate that Tawa’s aref abstraction and automatic warp specialization provide a robust orchestration mechanism that overlaps TMA loads and WGMMA across precisions.
@@ V-C GEMM Variants
To assess the generalizability of Tawa across diverse kernel patterns, we evaluate two GEMM variants implemented in Triton. The first is batched GEMM, which executes multiple small GEMMs of identical shape within a single kernel. The second is grouped GEMM, which processes multiple GEMMs of varying shapes in one kernel. Both patterns are pervasive in Mixture of Experts models [40, 41, 42]. For batched GEMM, we fix the batch size to 8 and sweep , , and across a range of values. For grouped GEMM, we vary the number and shapes of GEMMs. As shown in Fig. 9, Tawa consistently outperforms Triton baselines in both cases, achieving speedups of up to 7%. The improvements arise from Tawa’s aref-based partitioning and automatic warp specialization, which allow data movement for one GEMM to be overlapped with computation from another. In batched GEMM, this overlap reduces idle time between consecutive small multiplications, while in grouped GEMM, it schedules heterogeneous shapes more efficiently within a single kernel.
@@ V-D Multi-Head Attention
For FP16, Tawa attains up to 96% of FA3 throughput and delivers a consistent 1.21 speedup over Triton. This gap largely stems from the Triton baseline being effectively a FlashAttention-2 style implementation [43], which does not fully exploit Hopper warp specialized producer and consumer pipelines, in particular TMA driven global to shared transfers overlapped with WGMMA on Tensor Cores. At short sequences, the advantage of warp specialization is muted because prologue, epilogue, and barrier costs are not yet amortized. As grows to at least 4K, memory traffic and synchronization become dominant; in this regime Tawa outperforms TileLang and ThunderKittens by up to 1.13 and 1.23, respectively, indicating more effective overlap between data movement and compute. Although causal MHA reduces effective Tensor Core utilization due to mask induced control and data hazards, Tawa sustains strong performance at long sequences and reaches its peak at K.
For FP8, the benefits of automatic warp specialization are even more pronounced at large . Tawa reaches up to 89% of FA3 and achieves 1.11 speedup over Triton. TileLang and ThunderKittens failed to execute our FP8 attention configurations, suggesting that their kernels are tuned primarily for FP16 and lack the layout management and pipeline scheduling needed to handle FP8 WGMMA paths robustly.
@@ VI Future Work
Looking further ahead, we aim to generalize Tawa beyond Hopper to support next-generation architectures such as Blackwell. In particular, Blackwell introduces tensor memory (tmem) as a new hardware-managed memory hierarchy that complements TMA by providing more flexible data orchestration between shared and global memory. Leveraging tmem effectively will require extending aref to model multi-level data movement and revisiting scheduling decisions across heterogeneous memory resources. In addition, future workloads demand even greater scalability, motivating support for multiple producers and consumers for multiple warp groups, as well as more sophisticated graph partitioning algorithms that account for load balancing, register pressure, and memory footprint simultaneously. By incorporating these extensions, Tawa could evolve into a general-purpose compiler framework for warp-specialized, asynchronous GPU programming across successive hardware generations.
@@ VII Conclusion
In this paper, we presented Tawa, an automated compiler that generates efficient warp-specialized kernels for modern GPUs through the introduction of the asynchronous reference. Our evaluation demonstrates that Tawa attains performance comparable to highly optimized handwritten libraries across a range of benchmarks, while significantly reducing programming effort. We believe this work offers a step toward principled compiler support for asynchronous GPU programming and can inspire future research on automatically generating high-performance kernels for next-generation architectures.

### v2

@@ I Introduction
To address these challenges, we propose Tawa 111A shorthand for Task-Aware Warp Specialization with Aref. The implementation has been upstreamed to OpenAI Triton: https://github.com/triton-lang/triton/tree/aref_auto_ws, with the first PR available at #6288. an automated compiler that systematically analyzes computation graphs generated from a tile-based DSL, partitions programs into warp-specialized components, and manages both communication and pipelining. Our contributions are as follows:
We propose asynchronous references (aref), an abstraction for warp-level communication with formal guarantees. It captures the intent of asynchronous data movement without exposing hardware-specific intrinsics, enabling a clean and portable representation at the IR level.
We develop Tawa, the first fully automated compilation flow for warp specialization on modern NVIDIA GPUs. Tawa operates on unmodified, annotation-free Triton programs, performs task-aware partitioning across warp groups, applies multi-granularity pipelining adapted to diverse kernel structures, and emits high-performance PTX code for GPU execution.
@@ II-B Existing GPU Programming Models
To further raise the abstraction level, domain-specific compiler infrastructures have been proposed. TVM [14] pioneered an end-to-end compilation framework with automated scheduling [15, 16], while Relax [17] extends it with a graph-level IR for dynamic workloads. Fireiron [18] also leverages a schedule language for specifying the tiling and data movement, and Graphene [19] further enhances this idea by providing an IR for mapping tensors with various layouts to Tensor Core instructions. While powerful, these systems were developed prior to Hopper and do not directly exploit TMA and hardware warp specialization, leaving significant performance untapped. Similarly, lots of efforts on GPU warp specialization [20, 21, 22] were proposed before the Hopper architecture, which do not effectively use the latest TMA features. WASP [23] considers GPUs with TMA but requires architectural modifications and runs solely on a simulator. In contrast, Tawa targets commercial NVIDIA GPUs and demonstrates measurable speedups on real workloads.
Motivated by the need for finer control, a new class of tile-based kernel languages has emerged. Triton [8], built on MLIR [24], allows programmers to write efficient kernels in Python while abstracting tiling and memory management. Despite its success in democratizing kernel development, Triton, before this work, still lacks native abstractions for warp specialization and instead relies on Ampere-style software pipelining scheme with cp.async, leaving a performance gap on the Hopper architecture. Recent work introduces explicit APIs for dataflow scheduling and pipeline composition. Cypress[25] proposes a task-based programming model that abstracts complex data movement and synchronization through logical computation descriptions and mapping specifications. However, programmers are still required to manually manage the multi-level hierarchy of threads, warps, and SMs, along with the corresponding data decomposition. Gluon [26] proposes an IR for Triton that adds layout abstractions but still requires programmers to explicitly incorporate communication operations. These approaches provide expert users with expressive control, but they continue to impose significant responsibility for synchronization and resource management, which limits programmer productivity.
Additionally, TileLang [27, 28] is a new tile-based GPU kernel language with built-in warp specialization support. It implicitly schedules operations to form pipelines, but without a formal correctness guarantee and with limited control over fine-grained MMA pipelines. In contrast, Tawa formally defines asynchronous communication semantics via the aref abstraction (§III-B), treating communication as a first-class IR construct. This design ensures that the compiler generates correct warp-specialized code by construction, preventing deadlocks such as premature data access in aref objects. Tawa ’s multi-granularity pipelining further leverages aref to guarantee correctness across diverse communication patterns. Importantly, Tawa requires no user annotations or code modifications, whereas TileLang still requires marked pipelined loops (T.pipelined) and explicit copy operations (T.copy) to define pipeline stages.
Taken together, these trends highlight a growing need for raising the level of abstraction. Users should not cope with low-level hardware details or manually orchestrate warp-group execution. Tawa bridges these gaps with a fully automated compiler equipped with aref.
@@ III-A Overview
Fig. 2(a) illustrates the overall design of Tawa. Since Triton [8] has become the de facto backend for PyTorch [29, 30] and is widely adopted for writing efficient GPU kernels, we select Triton as the frontend interface of our compiler. Programmers write kernels in Triton-Python with tiled computation and TMA communication in a CTA (Fig. 2(b)), which are first translated into the standard Triton-MLIR representation. Building upon this infrastructure, Tawa introduces a series of compiler passes that automate the generation of warp-specialized programs directly from high-level Python code. Most importantly, we require no modifications to existing Triton code; users can simply set enable_warp_specialization=True when calling the Triton kernel, and it only takes effect on Hopper and newer GPU architectures.
@@ III-C1 Partition Annotation
We perform a backward traversal along the use-def chains starting at the kernel’s side-effecting sinks (e.g., Store). During the traversal, we attach a semantic tag to every node by inspecting its effects. Nodes that contribute to address computation, such as the pointer arithmetic for data transfers, are marked as iteration statements (orange edges). Nodes that transform or consume a tile for actual computation, such as WGMMA, are marked as tile statements (blue edges). These tags make the high-level intent of each region explicit. Because iteration statements are often scattered throughout the IR, semantic tagging is necessary to recover producer-related operations. For example, in Fig. 2(b), the address o_k calculation in L20 is separated with the TMALoad operation in L16-17.
After tagging, Tawa constructs the partitions. A partition is a dependency-closed subgraph that represents one warp role and will be executed by one warp group. To determine these subgraphs, Tawa identifies the cut that separates producer behavior from consumer behavior. All iteration statements, together with the TMALoad operations they dominate, are grouped into the producer partition. All tile statements and their dependents form the consumer partition. The graph is then closed under dependencies, so each partition contains all operations required for correctness. If a node is used by both partitions, for example an address calculation that contributes to both a load and a store, Tawa duplicates that computation. This guarantees that each partition is self-contained, and it avoids creating cross-partition data dependencies that would violate warp specialization constraints.
On Hopper, the typical configuration is two partitions. The producer partition corresponds to the load warp group (WG0), and the consumer partition corresponds to the compute warp group (WG1), as illustrated in Fig. 5(b). The epilogue is attached to WG1 so that output writes occur exactly once. For Blackwell and later architectures that provide more warp roles, Tawa can create additional partitions. The same semantic analysis drives these refinements, and the graph cut is adjusted to form separate load, compute, reduction, or epilogue subgraphs. In all cases, partitions are built from semantic roles, and duplication is applied when required so that each warp group receives a complete and independent subgraph for execution.
@@ III-C2 Loop Distribution
After the operations are annotated, we need to construct the actual IR that divides the workload for each WG. For each cross-partition edge (e.g., TMALoadLocalAlloc in Fig. 5(a)), we create an aref tensor with size , denoting a -slot cyclic buffer. is chosen to maximize overlap between TMA transfers and Tensor Core computation. A larger allows the producer to run further ahead and hide more TMA latency, while a smaller reduces buffer footprint and synchronization overhead. We study the effect of in § V-E. A put operation is created on the producer side that places each TMALoad result into an aref slot. The slot index is computed as the iteration modulo the buffer size , allowing the buffer slots to be reused (L15, L23 in Fig. 2(c)). Similarly, a get operation is created on the consumer side that gets the data from the same aref slot. After the data are used for computation in WGMMA, an additional consumed operation is inserted at the end to indicate the slot has been freed. As mentioned in § III-B, aref supports grouping multiple values in the buffer to share the mbarrier. Therefore, a straightforward optimization when creating aref is to check whether the loaded tensors are used in the same WGMMA. If so, we can put a tuple of tensors in the aref buffer and share the channel for communication between WGs. The resulting computation graph is shown in Fig. 5(b). Since aref has already carried buffers, the LocalAlloc operations can also be eliminated but to directly read from the aref buffer.
@@ V-A Experimental Setup
Tawa is built on top of Triton [8] (commit: 0c7edf) with about 4K lines of C++ and introduces a compact MLIR dialect that encapsulates aref operations. We compare Tawa against a set of state-of-the-art GPU programming frameworks and libraries, including the closed-source commercial library cuBLAS [39] (v12.7), the NVIDIA open-source CUDA linear algebra library CUTLASS [9] (v4.0.0), and the baseline Triton under the same commit. We also include two academic frameworks: ThunderKittens[10] (6c27e2), a CUDA-based AI kernel library, and TileLang[27] (422fb1), a TVM [14]-based DSL. To ensure a fair comparison, all baselines are hand-tuned with a fixed set of tile sizes (combinations of ). In the default Triton implementation, the official software pipelining is enabled, with pipeline depth predetermined to balance overlap between communication and computation. For kernels generated by Tawa, the size of the aref and the depth of the MMA pipeline are selected manually to maximize performance, providing each approach with equivalent opportunity to exploit the underlying hardware.
@@ V-B Matrix Multiplication
We evaluate GEMM kernels with matrix dimensions of while sweeping the inner dimension from to , and report the throughput for FP16 and FP8. Notice these matrix sizes are representative in real LLM workloads [40, 41, 42, 43]. As shown in Fig. 8, across both precisions, Tawa achieves the same level of performance as the highly optimized kernel library cuBLAS while outperforming the other general-purpose frameworks on most shapes. Tawa reaches up to 79% hardware utilization in both FP16 and FP8.
In FP16, Tawa delivers average speedups of 1.01, 1.13, 1.15, and 1.09 over cuBLAS, Triton, TileLang, and ThunderKittens, respectively. The improvements over Triton highlight the advantage of overlapping TMA-driven data movement with WGMMA execution. In contrast, Triton employs an Ampere-style software pipelining scheme for asynchronous copies, which is less effective on Hopper than the hardware-backed warp-specialization strategy in Tawa. We also observe Tawa is worse than cuBLAS for small , which is mainly because the overhead of Triton becomes relatively significant compared to the highly optimized CUDA library implementation in cuBLAS. TileLang and ThunderKittens are extensively tuned for large , leading to stronger performance than Tawa when .
In FP8, the benefits of automatic warp specialization are more pronounced, particularly at small . On average, Tawa achieves 1.06, 1.02, 1.22, and 1.24 speedup over cuBLAS, Triton, TileLang, and ThunderKittens, respectively. This is because the smaller, faster FP8 tiles make Tensor Core computation so rapid that memory transfer and synchronization become bottlenecks. Tawa’s warp-specialized pipeline alleviates this by maintaining continuous data flow. The aref abstraction automatically balances producer and consumer warps and deepens prefetching to fully utilize Tensor Cores, yielding larger relative gains for FP8 than for FP16. TileLang and ThunderKittens appear less tuned for FP8, resulting in up to 1.59 and 1.61 lower throughput at small , respectively. Overall, these results indicate that Tawa’s aref abstraction and automatic warp specialization provide a robust orchestration mechanism that overlaps TMA loads and WGMMA across precisions. We also observe that the Triton baseline exceeds Tawa in certain cases, primarily because Triton has extensively tuned its software pipeline for simple GEMM kernels. Tawa’s advantages become more pronounced in more complex kernels with deeper communication–computation dependencies, such as FlashAttention in § V-D.
@@ V-C GEMM Variants
To assess the generalizability of Tawa across diverse kernel patterns, we evaluate two GEMM variants. The first is batched GEMM, which executes multiple small GEMMs of identical shape within a single kernel. The second is grouped GEMM, which processes multiple GEMMs of varying shapes. Both patterns are pervasive in Mixture of Experts models [44, 45, 42]. We compare against Triton and TileLang; ThunderKittens does not provide functioning kernels for these cases. For batched GEMM, we fix the batch size to 8 and sweep , , and from 1K to 16K with square inputs. For grouped GEMM, we vary both the number and shapes of GEMMs, fixing and while letting be a multiple of 512. As shown in Fig. 9, Tawa consistently outperforms the Triton baseline in both settings, achieving speedups of up to 7%. Compared to TileLang, Tawa is up to 50% faster on batched GEMM. In grouped GEMM, TileLang performs well for small groups but degrades as group size increases. These gains stem from Tawa’s aref-based partitioning and automatic warp specialization, which allow data movement for one GEMM to be overlapped with computation from another. In batched GEMM, this overlap reduces idle time between consecutive small multiplications, while in grouped GEMM, it schedules heterogeneous shapes more efficiently within a single kernel.
@@ V-D Multi-Head Attention
For FP16, Tawa attains up to 96% of FA3 throughput and delivers a consistent 1.21 speedup over Triton. This gap largely stems from the Triton baseline being effectively a FlashAttention-2 style implementation [46], which does not fully exploit Hopper warp specialized producer and consumer pipelines, in particular TMA driven global to shared transfers overlapped with WGMMA on Tensor Cores. At short sequences, the advantage of warp specialization is muted because prologue, epilogue, and barrier costs are not yet amortized. As grows to at least 4K, memory traffic and synchronization become dominant; in this regime Tawa outperforms TileLang and ThunderKittens by up to 1.10 and 1.23, respectively, indicating more effective overlap between data movement and compute. Although causal MHA reduces effective Tensor Core utilization due to mask induced control and data hazards, Tawa sustains strong performance at long sequences and reaches its peak at K.
For FP8, the benefits of automatic warp specialization are even more pronounced at large . Tawa reaches up to 89% of FA3 and delivers 1.11 and 1.48 speedup over Triton and TileLang. ThunderKittens fails to run our FP8 attention configurations, indicating that its kernels are primarily tuned for FP16 and lack the layout management and pipeline scheduling needed for efficient FP8 execution.
@@ VI Future Work
Looking further ahead, we aim to generalize Tawa beyond Hopper to support next-generation architectures such as Blackwell. In particular, Blackwell introduces tensor memory (tmem) as a new hardware-managed memory hierarchy that complements TMA by providing more flexible data orchestration between shared and global memory. Leveraging tmem effectively will require extending aref to model multi-level data movement and revisiting scheduling decisions across heterogeneous memory resources. In addition, future workloads demand even greater scalability, motivating support for multiple producers and consumers for multiple warp groups with distinct roles, as well as more sophisticated graph partitioning algorithms that account for load balancing, register pressure, and memory footprint simultaneously. Specifically, these fine-grained roles can be annotated during the task-aware partitioning process based on instruction semantics as discussed in § III-C. For example, designating specialized functions such as softmax as special-function warps, while computations outside the main loop form an epilogue warp group that utilizes tmem.
@@ VII Conclusion
In this paper, we introduced Tawa, an automated compiler that generates efficient warp-specialized kernels for modern GPUs using asynchronous references. Our evaluation shows that Tawa delivers performance on par with highly optimized handwritten libraries while greatly reducing programming effort. We believe this work offers a step toward principled compiler support for asynchronous GPU programming and can inspire future research on automatically generating high-performance kernels for next-generation architectures.
Acknowledgement
We would like to thank Zihao Ye and the OpenAI Triton team for valuable feedback during the initial design of Tawa. We also thank Rohan Yadav, Matthias Springer, the anonymous reviewers, and our shepherd for their insightful comments on earlier drafts of this manuscript. Finally, we thank Lei Wang for providing guidance on setting up the TileLang environment. Hongzheng Chen and Zhiru Zhang are supported in part by ACE, one of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA and NSF Award #2118709.